# Quantum Learning and Estimation for Distribution Networks and Energy Communities Coordination

### v1

@@ III-B Quantum Amplitude Estimation
Thus, the probability of measuring the ancillary qubit in state $|1\rangle$ is $\sin^{2}((2k+1)\theta)$. When $k$ is chosen to be $\lfloor\frac{\pi}{4\theta}\rfloor$, we have $\sin^{2}((2k+1)\theta)\approx 1$. Here, we achieve the amplitude amplification of the target state $|1\rangle$. It has been verified that the estimated value $\tilde{a}$ satisfies $|\tilde{a}-a|\leq O(\frac{1}{M})$ with probability of at least $\frac{8}{\pi^{2}}$. $M$ is the number of oracle queries. QAE algorithm has been verified to be able to achieve a quadratic speedup compared to the $O(\frac{1}{\sqrt{M}})$ convergence rate of classical MC methods [22, 23], where $M$ is the number of MC simulations.
By applying QAE on (4), (4) can be computed efficiently.
@@ IV-A Settings
A modified IEEE 33-bus distribution network integrated with DGs, uncertain loads, and multi-type ECs is adopted as the test system. One WT and two PVs are connected to buses 5, 11, and 31, respectively. Uncertain loads are located at buses 10, 12, and 14, while the remaining loads are assumed to be deterministic. Three industrial, commercial, and residential EC are connected to buses 7, 15, and 18, respectively. The voltage magnitude is restricted as $|V_{i}|\in[0.90,1.10]\ \text{(p.u.)},\ \forall i\in\varOmega^{\mathrm{N}}$. $\Delta t$ = 15 min and $T$ = 96. $\lambda$ = 1 and $\alpha$ = 0.95. The VQC for Q-TCN model uses 8 qubits and 3 layers. The VQC in Q-LSTM cell uses a hidden size of 3 and 2 recurrent layers. The learning rate is set as 0.01. Three datasets comprising 10,240 incentive-response samples from three ECs are generated using the models provided in Appendix A and utilized to train the proposed mapping model. Note that since modeling complex ECs’ response behaviors is not the focus of this work, the models are simplified in form. Nevertheless, the proposed quantum learning and estimation approach remains applicable to more complex cases. QNNs are implemented using Pennylane, Jax, and Flaxnn libraries. QAE is implemented using Qiskit library. All computations are implemented in Python 3.10.11, conducted on a Windows 11 64-bit operating system with an Intel Core i9-13900HX @ 2.30 GHz processor and 16 GB RAM.
IV-B Performance of Q-TCN-LSTM
In this part, we evaluate the performance of Q-TCN-LSTM in terms of accuracy and computational efficiency. The loss curves of the classical TCN-LSTM and Q-TCN-LSTM models on the test set during the training process are presented in Fig. 8. As observed, both models exhibit a rapid decrease in loss during the initial training epochs, followed by a gradual decline towards a stable convergence. This behavior highlights the effectiveness of the combined short-term and long-term temporal feature extraction capabilities of both the TCN–LSTM and Q-TCN–LSTM models. TCN-LSTM achieves convergence at around 25 epochs, with a final MSE loss of 0.013. In contrast, the Q-TCN-LSTM model demonstrates faster convergence at around 10 epochs and eventually reaches a relatively low MSE loss of 0.004. These results demonstrate the superior learning ability of the Q-TCN-LSTM model.
Figure 8: Loss curves of TCN-LSTM and Q-TCN-LSTM on commercial EC.
To further illustrate the effectiveness of the Q-TCN-LSTM model in capturing the incentive-response behavior of ECs, we visualize two randomly selected response trajectories of the commercial EC from the test set, as shown in Fig. 9. The mapping results (blue dotted line) closely match the true values (red solid line) with minimal mapping deviation, demonstrating the model’s high mapping precision and temporal generalization performance.
Figure 9: Mapping results of two samples of commercial EC.
@@ IV-B Performance of Q-TCN-LSTM
Figure 10: Mapping performance comparison of multiple models.
@@
As shown in Fig. 10, QNNs consistently outperform their classical counterparts across all three ECs, achieving lower loss and demonstrating superior function approximation capabilities. This advantage can be attributed to the unique properties of quantum computing. Specifically, quantum entanglement provides QNNs with powerful nonlinear expressive capabilities, enabling the effective modeling of complex relationships between pricing incentives and responses within ECs.
@@
Among the QNN architectures evaluated, Q-TCN-LSTM achieves the highest mapping accuracy. This is primarily due to its hybrid architecture, which integrates the strengths of both Q-TCN and Q-LSTM networks. In Q-TCN, the integration of quantum convolutional layers with quantum parallelism enables the model to efficiently capture fine-grained temporal dependencies within short time sequences. The Q-LSTM leverages quantum superposition to significantly increase its memory capacity. This allows the model to maintain strong correlations between distant data points, leading to a better understanding of long-range context. By leveraging the strengths of both Q-TCN and Q-LSTM, the Q-TCN-LSTM architecture is capable of capturing both short-term and long-term temporal dependencies inherent in the EC response patterns, thereby enhancing its overall performance.
@@
Moreover, we compare the computational efficiency of Q-TCN-LSTM with other models. Due to the current limitations in quantum hardware availability, all quantum operations are simulated on classical processors. We adopt the formula in [28] to approximate the quantum computational time: $T^{\mathrm{Q}}=T^{\mathrm{P}}+T^{\mathrm{G}}\cdot D+T^{\mathrm{M}}$. where $T^{\mathrm{P}}$ denotes state preparation time, $T^{\mathrm{G}}$ represents the average duration of a single quantum gate, $D$ is the maximal circuit depth, and $T^{\mathrm{M}}$ is the quantum measurement time. Empirical values from [28] suggest that $T^{\mathrm{P}}+T^{\mathrm{M}}=1\ \mu s$ and $T^{\mathrm{G}}=10\ ns$, which indicates that the time for state preparation and measurement takes much part of the total time. Thus, one execution time of the convolution operation in the VQC in Q-TCN takes around 1.1 $\mu s$, which is much less than that of classical convolution and activation operation (around 18 $\mu s$). This is because quantum parallelism enables performing multiple operations simultaneously, significantly boosting computational speed. In terms of model size, the TCN-LSTM model has a total of 37353 parameters, while the Q-TCN-LSTM model uses only 94 parameters, approximately only 0.25% of the TCN-LSTM model’s parameter count. Q-TCN-LSTM outperforms the classical TCN-LSTM while using far fewer parameters and operating at a much higher speed, demonstrating its high computational efficiency.
IV-C Performance of QAE
@@
To validate the computational efficiency of quantum estimation methods for problems involving a large number of scenarios, we compare the estimation error (expressed as a percentage) and the corresponding computation time under varying numbers of MC samples and QAE qubits. The detailed results are illustrated in Fig. 11.
Figure 11: Comparison of MC and QAE: (a) MC, (b) QAE with Circuit 1.
@@
As shown in Fig. 11, MC achieves a low estimation error of approximately $0.001\%$ with a substantial number of $10^{6}$ samples, while QAE attains a comparable error level of $0.0008\%$ with only 12 qubits. This demonstrates a substantial reduction in required computational resources. Moreover, QAE demonstrates a marked advantage in time efficiency. Under a similar error threshold (MC with $10^{6}$ samples vs. QAE with 12 qubits), MC requires approximately 1.47 s of computation time, while QAE completes the task in approximately 0.01 $\mu s$. This efficiency gain is primarily attributed to the quadratic speedup in query complexity offered by QAE. Furthermore, given the exponential growth of quantum circuit execution capacity with increasing qubits, the computational advantage of QAE is expected to become even more pronounced in large-scale estimation problems. Overall, these findings underscore the superiority of QAE in terms of both estimation accuracy and computational efficiency.
@@
In addition, the performance of QAE using the approximate circuit (Circuit-1) and the precise circuit (Circuit-2) is compared in terms of estimation error and computational time, as shown in Table I. The results indicate that Circuit-1 does not exhibit notable increase in estimation error compared to Circuit-2. This is primarily due to the linearity of the function $\hat{f}(\cdot)$ in the current estimation task. In such cases, the corresponding quantum rotation angles can be effectively approximated by linear interpolation. However, if $\hat{f}(\cdot)$ exhibits strong nonlinearity, the linear approximation of rotation angles might become inaccurate, leading to potentially higher estimation errors. In contrast, Circuit-2 can represent such functions with higher precision but at the cost of increased complexity: it requires $12n+$ 1 qubits and $2^{n}$ quantum controlled rotation gates, resulting in substantially higher execution time. Therefore, in practical applications, the choice between the two circuit designs should be guided by the functional characteristics of the estimation target and the required trade-off between accuracy and computational complexity.
TABLE I: Performance Comparison of Circuit-1 and Circuit-2
Circuit Metric Number of Qubits
5 6 7 8 9 10 11
Circuit-1 Error (%) 0.0939 0.0470 0.0236 0.0119 0.0060 0.0031 0.0016
Time ($\mu$s) 0.0040 0.0060 0.0060 0.0060 0.0060 0.0061 0.0063
Circuit-2 Error (%) 0.0937 0.0468 0.0234 0.0117 0.0058 0.0029 0.0014
Time ($\mu$s) 0.1685 0.8326 4.8751 93.423 755.83 6534.4 58568
IV-D Result of DNs-ECs Coordination
@@
In the optimization problem of DNs, the explicit gradient structure is established between the objective function and the decision variables, facilitating the search for the optimum. Thus, leveraging the accurate incentive-response mapping of ECs enabled by Q-TCN-LSTM, along with the accurate gradient estimation enabled by QAE, gradient descent can be efficiently employed. The bilevel DNs-ECs coordination optimization results are shown in Fig. 12.
@@
Figure 12: Results of price and trading strategies of ECs: (a) Electricity trading price, (b) Residential EC, (c) Commercial EC, (d) Industrial EC.
@@
As illustrated in Fig. 12, the optimized pricing scheme (Price 2) exhibits a more intricate and time-varying structure compared to the baseline fixed pricing (Price 1). This enables better alignment with the comprehensive load characteristics and operational profiles of multi-type ECs. Figure. 12 indicates that Price 2 increases during 0-8 h and 16-24 h, while decreasing significantly during 8-16 h. This pricing pattern closely matches the underlying net load dynamics of the ECs: during 8-16 h, high penetration of renewable generation leads to lower net loads, making it suitable to reduce electricity prices and encourage local consumption. Conversely, the 16-24 h period corresponds to the daily peak demand interval, where elevated prices are employed to suppress load aggregation and incentivize demand shifting. The effectiveness of this pricing scheme is further reflected in user response behavior, as shown in Fig. 12(b)-(d). Following price optimization, all three user types exhibit varying degrees of net load shifting, particularly during the high-price hours (16-24 h), resulting in a noticeable reduction in peak net loads. Among them, industrial users demonstrate the most substantial response due to their large and adjustable load capacity, indicating high sensitivity to price signals. Commercial users show moderate responsiveness, while residential users exhibit limited flexibility due to the inherent rigidity of their consumption patterns and fewer controllable loads. This coordinated response behavior effectively reduces the negative impact of ECs on the DNs. After optimizing the pricing strategy, the average penalty cost of voltage violation decreases from 17.76 to 2.24, indicating the capacity of user-side response to alleviate potential risks and enhance distribution network reliability. Overall, leveraging the advantages of quantum computing, the DNs-ECs coordination optimization problem can be solved efficiently.

### v2

@@ III-B1 General Methodology of QAE
Thus, the probability of measuring the ancillary qubit in state $|1\rangle$ is $\sin^{2}((2k+1)\theta)$. When $k$ is chosen to be $\lfloor\frac{\pi}{4\theta}\rfloor$, we have $\sin^{2}((2k+1)\theta)\approx 1$. Here, we achieve the amplitude amplification of the target state $|1\rangle$. It has been verified that the estimated value $\tilde{a}$ satisfies $|\tilde{a}-a|\leq O(\frac{1}{M^{\mathrm{Q}}})$ with probability of at least $\frac{8}{\pi^{2}}$. $M^{\mathrm{Q}}$ is the number of oracle queries. QAE algorithm has been verified to be able to achieve a quadratic speedup compared to the $O(\frac{1}{\sqrt{M^{\mathrm{S}}}})$ convergence rate of classical MC methods [25, 29] under relatively ideal conditions, where $M^{\mathrm{S}}$ is the number of MC simulation samples.
III-B2 Method of applying QAE to (4)
Next, we propose a quantum estimation method for efficient computation of (4). Based on the LinDistFlow model in (1), variables such as nodal voltage and line power flow exhibit a linear relationship with the uncertain variables (i.e., RES generation, load demand) and $P^{\mathrm{EX,EC}}$ [33, 34]. Thus, gradient terms $\partial(\pi^{\mathrm{EC}}_{t}P^{\mathrm{EX,EC}}_{i,t})/\partial\pi^{\mathrm{EC}}_{t}$ and $\partial P^{\mathrm{EX,DN}}_{s,t}/\partial\pi^{\mathrm{EC}}_{t}$ can be easily computed. Regarding the penalty terms with $\cdot$$_{+}$, note that the distributions of $V_{i,s,t}$ and $P_{ij,s,t}$ can be accurately determined based on the known distribution of uncertain variables, the relationship between these penalty gradient terms (e.g., $\frac{\partial[V_{i,s,t}-\overline{V}_{i}]_{+}}{\partial\pi^{\mathrm{EC}}_{t}}$) and scenario index $i$ presents a piecewise characteristic. Note that the computational procedures for these penalty terms are essentially the same. Take one penalty term $\sum_{s=1}^{S}\gamma_{s}\frac{\partial[V_{i,s,t}-\overline{V}_{i}]_{+}}{\partial\pi^{\mathrm{EC}}_{t}}$ as an example, the core steps are as follows: a) encoding the distribution of $V_{i,t}$ using a quantum state preparation circuit, b) adopting specific controlled rotation circuits to model the penalty term, where Circuit-1 in Fig. 6 can be utilized for implementation if the linear approximation error is acceptable, otherwise Circuit-2 in Fig. 7 can be employed if the original $f(i)$ is retained. Such a choice can be determined based on the trade-off between accuracy and computational complexity. c) applying a QAE block to the auxiliary qubit to obtain the estimation result. Similarly, other penalty terms can also be computed efficiently. Based on the above analysis, (4) can be computed efficiently by applying QAE.
@@ IV-A Settings
A modified IEEE 33-bus distribution network integrated with DGs, uncertain loads, and multiple ECs is adopted as the test system. One WT and two PVs are connected to buses 5, 11, and 31, respectively. Three uncertain loads are located at buses 10, 12, and 14, while the remaining loads are assumed to be deterministic. An industrial EC, a commercial EC, and a residential EC are connected to buses 7, 15, and 18, respectively. The voltage magnitude is restricted as $|V_{i}|\in[0.90,1.10]\ \text{(p.u.)},\ \forall i\in\varOmega^{\mathrm{N}}$. $\Delta t$ = 15 min and $T$ = 96. $\lambda$ = 1 and $\alpha$ = 0.95. The VQC for Q-TCN model uses 4 qubits and 3 layers. The VQC in the Q-LSTM cell uses a hidden size of 3 and 2 recurrent layers. The learning rate is set to 0.01. Three datasets comprising 10,240 incentive-response samples from three ECs are utilized to train the proposed mapping model. QNNs are implemented using PennyLane, Jax, and Flaxnn libraries. QAE is implemented using Qiskit library. All computations are implemented in Python 3.10.11, conducted on a Windows 11 64-bit operating system with an Intel Core i9-13900HX @ 2.30 GHz processor and 16 GB RAM.
IV-B Performance of Quantum Learning
We evaluate the performance of Q-TCN-LSTM in terms of accuracy and computational efficiency. In terms of convergence speed, compared to the classical TCN-LSTM, which converges in about 25 epochs with an MSE of 0.013, Q-TCN-LSTM converges faster within 10 epochs and achieves a lower MSE of 0.004. These results demonstrate the improved learning ability of Q-TCN-LSTM, benefiting from quantum advantage. Furthermore, we visualize two randomly selected response trajectories of the commercial EC from the test set, as shown in Fig. 8. The mapping results (blue dotted line) closely match the true values (red solid line) across both high-value and low-value ranges with minimal mapping deviation, demonstrating the model’s high mapping precision and temporal generalization performance.
Figure 8: Mapping results for two commercial EC samples.
@@ IV-B Performance of Quantum Learning
Figure 9: Performance comparison of multiple models.
@@
As shown in Fig. 9, QNNs consistently outperform their classical counterparts across all three ECs, achieving lower MSE and demonstrating superior function approximation capabilities. This advantage can be attributed to the unique properties of quantum computing. Specifically, quantum entanglement provides QNNs with powerful nonlinear expressive capabilities, enabling the effective modeling of complex relationships between pricing incentives and ECs’ responses.
@@
Among the QNN architectures evaluated, Q-TCN-LSTM achieves the highest mapping accuracy. This is primarily due to its hybrid architecture, which integrates the strengths of both Q-TCN and Q-LSTM networks. In Q-TCN, the integration of quantum convolutional layers enables the model to efficiently capture fine-grained temporal dependencies within short time sequences. The Q-LSTM leverages quantum superposition to significantly increase its memory capacity. This allows the model to maintain strong correlations between distant data points, leading to a better understanding of long-range context. By leveraging the strengths of both Q-TCN and Q-LSTM, the Q-TCN-LSTM architecture is capable of capturing both short-term and long-term temporal dependencies inherent in the EC response patterns, thereby enhancing its overall performance.
@@
Moreover, we compare the computational efficiency of Q-TCN-LSTM with other models. Due to the current limitations in quantum hardware availability, all quantum operations are simulated on classical processors. Thus, for the following time efficiency computation, we separately measure two types of time indicators: simulation runtime ($\tau^{\mathrm{s}}$), which is the running time of task execution based on a quantum simulator on a classical computer; and quantum runtime ($\tau^{\mathrm{q}}$) which is the estimated time for an ideal quantum hardware device to run the corresponding task. Specifically, we adopt the formula in [35] to approximate the quantum runtime: $T^{\mathrm{Q}}\approx T^{\mathrm{P}}+T^{\mathrm{G}}\cdot D+T^{\mathrm{M}}$, where $T^{\mathrm{P}}$ denotes state preparation time, $T^{\mathrm{G}}$ represents the average duration of a single quantum gate, $D$ is the maximal circuit depth, and $T^{\mathrm{M}}$ is the quantum measurement time. Empirical values from [35] suggest that $T^{\mathrm{P}}+T^{\mathrm{M}}=1\ \mu s$ and $T^{\mathrm{G}}=10\ ns$, which indicates that the time for state preparation and measurement accounts for a significant portion of the total time  [^2] Note that the runtime of quantum algorithms on actual hardware is affected by multiple factors, such as the specific qubit platform, measurement methods, and error correction mechanisms, among others. Specific quantum runtime measurements are also reported in [36, 37, 38]. . Thus, one execution of the convolution operation in the VQC in Q-TCN takes around 0.03 $s$ of simulation runtime and 1.1 $\mu s$ of quantum runtime, which is much less than that of classical convolution and activation operation (around 0.017 $s$). This is because quantum parallelism enables performing multiple operations simultaneously, significantly boosting computational speed. In terms of model size, the TCN-LSTM model has a total of 37353 parameters, while the Q-TCN-LSTM model uses only 94 parameters, approximately only 0.25% of the TCN-LSTM model’s parameter count. Q-TCN-LSTM outperforms the classical TCN-LSTM while using far fewer parameters and operating at a much higher speed, demonstrating its high computational efficiency.
To investigate the impacts of noise on model performance, which is inherent in practical quantum hardware in the NISQ era, we simulated multiple noise levels during the model training process (theoretically [0, 1], with practical simulations focusing on [0, 0.2]). The relationship between the mapping performance (measured by MSE) and the noise level is:
TABLE I: Model Performance under Different Noise Levels
Noise Level 0 0.05 0.1 0.15 0.2
MSE 0.004 0.005 0.006 0.01 0.12
This result indicates that with the increase in noise level, the MSE shows a gradual upward trend. The core reason is that noise in quantum devices interferes with the stability of quantum states, leading to errors in the quantum computing process and thus reducing the model mapping accuracy. Therefore, in practical quantum hardware, quantum error correction strategies are needed to suppress noise interference and ensure computing performance.
IV-C Performance of Quantum Estimation
@@
To validate the computational efficiency of quantum estimation methods, we compare the estimation error (expressed as a percentage) and the corresponding computation time including both simulation runtime and quantum runtime under varying numbers of MC samples and QAE qubits. The detailed results are illustrated in Fig. 10.
Figure 10: Comparison of MC and QAE: (a) MC, (b) QAE with Circuit-2.
@@
As shown in Fig. 10, MC achieves a low estimation error of approximately $0.1\%$ with a substantial sample size of $10^{6}$, while QAE attains a comparable error level of $0.09\%$ with only 7 qubits. This demonstrates a substantial reduction in required computational resources. Moreover, QAE demonstrates a marked advantage in time efficiency. Under a similar error threshold (MC with $10^{6}$ samples vs. QAE with 7 qubits), MC requires approximately 1.41 s of computation time, while QAE completes the task in approximately 8.74 s of simulation runtime and around 920 $\mu s$ of quantum runtime. This efficiency gain is primarily attributed to the quadratic speedup in query complexity offered by QAE. Furthermore, given the exponential growth of quantum circuit execution capacity with increasing qubits, the computational advantage of QAE is expected to become even more pronounced in large-scale estimation problems. Overall, these findings underscore the superiority of QAE in terms of both estimation accuracy and computational efficiency.
@@
In addition, the performance of QAE using the approximate circuit (Circuit-1) and the precise circuit (Circuit-2) is compared in terms of estimation error and computational time, as shown in Table II. The results indicate that Circuit-1 exhibits a notable increase in estimation error compared to Circuit-2. This is primarily due to the linear approximation error. In contrast, Circuit-2 can represent such functions with higher precision but at the cost of increased complexity: it requires $2n+1$ qubits and $2^{n}$ quantum controlled rotation gates, resulting in substantially higher execution time. Therefore, in practical applications, the choice between the two circuit designs should be guided by the functional characteristics of the estimation target and the required trade-off between accuracy and computational complexity.
TABLE II: Performance Comparison of Circuit-1 and Circuit-2
Circuit Metric $n$
5 6 7 8 9 10
Circuit-1 $\epsilon$ (%) 9.123 6.688 5.459 4.834 4.534 4.379
$\tau^{\mathrm{s}}$ (s) 0.088 0.210 0.617 1.607 3.876 9.741
[…]
Circuit-2 $\epsilon$ (%) 2.933 2.534 0.093 0.042 0.017 0.008
$\tau^{\mathrm{s}}$ (s) 0.225 1.714 8.742 73.53 553.7 3534.4
[…]
IV-D Result of DNs-ECs Coordinated Operation
@@
In the optimization problem of DNs, leveraging the accurate incentive-response mapping of ECs enabled by Q-TCN-LSTM, along with the accurate gradient estimation enabled by QAE, gradient descent can be efficiently employed. The bilevel DNs-ECs coordinated operation optimization results are shown in Fig. 11.
@@
Figure 11: Results of incentive price and responses of ECs: (a) Electricity trading price, (b) Residential EC, (c) Commercial EC, (d) Industrial EC.
@@
As illustrated in Fig. 11, the optimized pricing scheme (Price 2) exhibits more time-varying characteristics compared to the baseline fixed pricing (Price 1). This enables better alignment with the comprehensive operation patterns of multiple ECs. Fig. 11 indicates that Price 2, compared with Price 1, exhibits a relatively higher price during hours 16-20, while presenting a relatively lower price during hours 8-16. This pricing pattern closely matches the underlying net load dynamics of the ECs: during hours 8-16, high penetration of renewable generation leads to lower net loads, making it suitable to reduce electricity prices to encourage load shifting. Conversely, during hours 16-20, the renewable generation decreases, and the load demand of ECs slightly increases, and thus the net load demand reaches a relative peak demand interval, where elevated prices are employed to suppress load aggregation and incentivize shifting the load from this period to other periods. The effectiveness of this pricing scheme is further reflected in user response behavior, as shown in Fig. 11(b)-(d). Following price optimization, all three ECs exhibit varying degrees of net load shifting, particularly during the high-price hours 16-20, resulting in a noticeable reduction in peak net loads. Among them, industrial users demonstrate the most substantial response due to their large and adjustable load capacity, indicating high sensitivity to price signals. Commercial users show moderate responsiveness, while residential users exhibit limited flexibility due to the inherent rigidity of their consumption patterns and fewer controllable loads. This coordinated response behavior effectively reduces the negative impact of ECs on the DNs. After optimizing the pricing strategy, the average penalty cost of voltage violation decreases from 50.9 to 21.8, indicating the capacity of user-side response to alleviate potential risks and enhance distribution network reliability. Overall, leveraging the advantages of quantum computing, the DNs-ECs coordinated operation optimization problem can be solved efficiently.